# Improving the Efficiency of Self-Supervised Adversarial Training through Latent Clustering-Based Selection

### v1

@@ D.1. Other Hyperparameter Analysis Details and Results
Table 4. Comparison results of self-supervised adversarial training algorithms with varying data selection schemes on SVHN. Instead of pseudo labels, the selected unlabeled data are assigned with the ground-truth SVHN labels provided in the dataset.
[…]
@@
D.2. SVHN with Ground-Truth Labels
We conduct additional experiments to test the performance of our data selection scheme on SVHN with pseudo labels replaced by ground-truth labels for the extra data. Table 4 shows the comparison results. With ground-truth labels, our LCS schemes can achieve similar robust accuracy by selecting just $10\%$ extra data compared with no selection result.
D.3. Pre-Trained CLIP as Intermediate Model
@@
In all our previous experiments, we began by training an intermediate model $f_{\hat{\theta}}$ using the same architecture as the final model $\theta_{\mathrm{final}}$. This intermediate model was trained exclusively on the labeled dataset and then used to generate pseudo labels and select a subset from the extra dataset. However, when dealing with generated or external data that already comes with labels, training an intermediate model specifically for selection undermines the selection process’s intended benefits, primarily to reduce computational and time complexity. In addition, in application scenarios where the size of the labeled dataset $\mathcal{S}_{l}$ is also limited, likely, the intermediate model can not be trained with high accuracy, leading to large amounts of inaccurate pseudo labels.
@@
To address these issues, we study the feasibility of using an alternative pre-trained model in LCS to perform data selection, specifically leveraging the CLIP model for extracting image embeddings (Radford et al., 2021). CLIP, trained on a vast array of internet images and text, offers a rich and versatile feature space that can be utilized without additional training. By using the capabilities of pre-trained CLIP, we can bypass the intermediate step of training a model from scratch, thereby streamlining the selection process and further enhancing overall efficiency. This substitution allows us to select a relevant subset without having to train an intermediate model from scratch. Our experiments indicate that using CLIP for data selection yields results that are comparable to those obtained using the trained intermediate model, both in terms of standard and robust accuracies. Figure LABEL:fig:comparison_clip shows our result when using CLIP instead of the previous intermediate model for data selection.
D.4. Resnet-18 Model Architecture
@@
Our main experiments focus on the WidResNet architecture. We further evaluate the performance of SSAT under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ using different data selection schemes and varying selection ratios with external unlabeled data on CIFAR-10 using a ResNet-18 model. Similar to our previous experiments, we train the ResNet-18 model on $50$K labeled samples from CIFAR-10 and a chosen subset of data from $500$K unlabeled images of Tiny-ImageNet. Our evaluation results are shown in Table 5, which suggests that our proposed LCS methods can greatly improve the efficacy of SSAT on ResNet-18 while achieving comparable robustness performance of “No Selection” with $\alpha=100\%$.
@@
Table 5. Comparisons of SSAT performance on CIFAR-10 using the ResNet-18 architecture under various configurations. The model training dataset consists of $50$K labeled CIFAR-10 images and $500$K unlabeled images drawn from Tiny-ImageNet.
[…]

### v2

@@ Appendix C Additional Experiments
TABLE IV: Comparison results of semi-supervised adversarial training algorithms with varying data selection schemes on SVHN. Instead of pseudo labels, the selected unlabeled data are assigned the ground-truth SVHN labels provided in the dataset.
$\textbf{Ratio }\bm{\alpha}$ Method Clean Accuracy ($\%$) Robust Accuracy ($\%$) Number of Epochs
@@
C-A Hyperparameter Sensitivity Analysis Details
In Section V-C, we present a detailed analysis of how accuracies are influenced by different hyperparameter values. In these experiments, we varied one hyperparameter at a time while keeping the others constant. Figure 4a illustrates how robust accuracy on the CIFAR-10 dataset changes with $\beta$ values when $20\%$ additional data is incorporated from the 11M data samples generated by the DDPM model, using a per-batch ratio of $0.3$. Figure 4b shows the variation in robust accuracy for the SVHN dataset as $\beta$ changes, with $10\%$ drawn from the SVHN extra dataset and a per-batch ratio of $0.5$.
In Figure 7a, we present how SSAT’s robust accuracy is affected on CIFAR-10 by varying $\gamma$ when $20\%$ extra data from the 11M generated DDPM samples is used, with $\beta$ fixed at $0.4$. Similarly, Figure 7b displays the effect of varying the per-batch ratio on robust accuracy for the SVHN dataset, using $10\%$ data from the SVHN extra dataset and fixing $\beta$ at $0.6$. In addition, Figures 4c and 4d explore how robust accuracy varies with the amount of data used. For CIFAR-10 experiments in Figure 4c, the per-batch ratio is set to $0.5$, and $\beta$ is fixed at $0.4$. For SVHN in Figure 4d, the per-batch ratio is set to $0.5$, and $\beta$ is set to $0.6$. We aim to capture the robust accuracy at the optimal epoch in each case, and we observe that for Figures 4c and 4d, as the amount of data increases, the optimal accuracy is reached at a considerably later epoch.
(a) CIFAR-10
(b) SVHN
Figure 8: Comparisons of SSAT methods with varying amounts of unlabeled data selected using LCS-KM with the CLIP model on (a) CIFAR-10 and (b) SVHN. We report standard and robust accuracies on the labeled test dataset.
C-B Pre-Trained CLIP as Intermediate Model
@@
In our experiments, we always began by training an intermediate model $f_{\hat{\theta}}$ using the same architecture as the final model $\theta_{\mathrm{final}}$. $f_{\hat{\theta}}$ was trained exclusively on the labeled dataset and then used to generate pseudo labels and select a subset from the extra dataset. However, when dealing with a generated or external dataset that already comes with labels, training an intermediate model specifically for selection undermines the data selection process’s intended benefits, primarily to reduce computational and time complexity. In application scenarios where the size of the labeled dataset $\mathcal{S}_{l}$ is also limited, likely, the intermediate model can not be trained with high accuracy, leading to large amounts of inaccurate pseudo labels.
@@
To address these issues, we study the feasibility of using an alternative pre-trained model in LCS to perform data selection, specifically leveraging the CLIP model for extracting image embeddings [52]. CLIP, trained on a vast array of internet images and text, offers a rich and versatile feature space that can be utilized without additional training. By using the capabilities of pre-trained CLIP, we can bypass the intermediate step of training a model from scratch, thereby streamlining the selection process and further enhancing overall efficiency. This substitution allows us to select a relevant subset without having to train an intermediate model from scratch. Our experiments indicate that using CLIP for data selection yields results that are comparable to those obtained using the trained intermediate model, both in terms of standard and robust accuracies. Figure 8 shows our result when using CLIP instead of the previous intermediate model for data selection.
C-C SVHN with Ground-Truth Labels
We conduct additional experiments to test the performance of our data selection scheme on SVHN with pseudo labels replaced by ground-truth labels for the extra data. Table IV shows the comparison results. With ground-truth labels, our LCS schemes can achieve similar robust accuracy by selecting just $10\%$ extra data compared with no selection result.
C-D Resnet-18 Model Architecture
@@
Our main experiments focus on the WideResNet architecture. We further evaluate the performance of SSAT under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ using different data selection schemes and varying selection ratios with external unlabeled data on CIFAR-10 using a ResNet-18 model. Similar to our previous experiments, we train the ResNet-18 model on 5050K labeled samples from CIFAR-10 and a chosen subset of data from 500500K unlabeled images of Tiny-ImageNet. Our evaluation results are shown in Table V, which suggests that our proposed LCS methods can greatly improve the efficacy of SSAT on ResNet-18 while achieving comparable robustness performance of “No Selection” with $\alpha=100\%$.
@@
TABLE V: Comparisons of SSAT performance on CIFAR-10 using the ResNet-18 architecture under various configurations. The model training dataset consists of 5050K labeled CIFAR-10 images and 500500K unlabeled images drawn from Tiny-ImageNet.
$\textbf{Ratio }\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$)
@@
Appendix D Extended Analyses
In this section, we further evaluate the flexibility of our method across various adversarial training algorithms and perturbation configurations (Section D-A), and examine the impact of intermediate model quality, particularly under the regime with insufficient labeled data (Section D-B). Without explicit mention, we use our LCS-KM method for selecting critical unlabeled data, as it derives the best-performing model when integrated with SSAT as suggested by Table I.
(a) TRADES ($\epsilon=4/255$, PGD-10)
(b) TRADES ($\epsilon=16/255$, PGD-10)
(c) AT ($\epsilon=8/255$, PGD-5)
(d) AT ($\epsilon=8/255$, PGD-10)
Figure 9: Illustration of SSAT performance on CIFAR-10 using LCS-KM under $\ell_{\infty}$ perturbations under various configurations: TRADES using $10$ PGD steps with (a) $\epsilon=4/255$ and (b) $\epsilon=16/255$; AT with $\epsilon=8/255$ using (c) 55 PGD steps and (d) $10$ PGD steps.
TABLE VI: Performance of intermediate models trained using different schemes, including self-supervised learning (SSL), pre-trained CLIP, and fully-supervised methods, on CIFAR-10 under varying ratios of labeled data. We also present the corresponding performance of the final model produced by SSAT with LCS-KM, where the ratio of selected unlabeled data $\alpha$ is set at $10\%$.
Labeled Ratio Intermediate Model Final Model (LCS-KM, $\alpha=10\%$)
Training Method Clean Acc ($\%$) Clean Acc ($\%$) Robust Acc ($\%$)
$1\%$ SSL ($5\%$ unlabeled) $78.8$ $82.9$ $39.8$
pre-trained CLIP $86.8$ $83.9$ $40.5$
$10\%$ SSL ($5\%$ unlabeled) $82.9$ $85.0$ $42.4$
pre-trained CLIP $86.8$ $85.3$ $43.2$
$20\%$ SSL ($5\%$ unlabeled) $87.0$ $87.2$ $48.3$
pre-trained CLIP $86.8$ $85.6$ $47.7$
100% fully-supervised $90.5$ $86.2$ $58.2$
pre-trained CLIP $86.8$ $86.5$ $57.4$
D-A Further Exploration of Generalizability
We compare four different settings of SSAT on CIFAR-10: random selection with $\alpha$ ratio of unlabeled data, selecting $\alpha$ ratio using LCS-KM, using only labeled data ($\alpha=0\%$), and using the entire dataset ($\alpha=100\%$).
Perturbation Size. We consider external data from TinyImages as the unlabeled dataset. We investigate the effect of varying the $\ell_{\infty}$ perturbation size $\epsilon$, setting it to $4/255$ and $16/255$ while keeping TRADES as the adversarial training algorithm similar to the experiments in Table I. The results are depicted in Figures 9a and 9b, suggesting that our LCS-KM method is highly effective compared with random selection and no selection, almost approaching the robust accuracies achieved by full utilization of unlabeled data. The trends are shown to be consistent across varying perturbation sizes.
Adversarial Training. We explore the impact of replacing TRADES with vanilla AT as the training algorithm. Similarly, we use external data from TinyImages as the unlabeled dataset. The results for PGD-based AT are shown in Figure 9d. Additionally, we reduce the number of PGD steps from $10$ to $5$, and the outcomes are summarized in Figure 9c. Across all configurations, our LCS-KM selection consistently outperforms random selection and achieves accuracy comparable to using the entire labeled dataset, suggesting the generalizability and robustness of our methods.
(a) External data
(b) Generated data
Figure 10: Illustration of SSAT performance on CIFAR-10 under $\ell_{2}$ perturbations with $\epsilon=128/255$ using LCS-KM by varying $\alpha$: (a) external TinyImages and (b) DDPM-generated data.
$\bm{\ell}_{2}$ Perturbations. We further evaluate the performance of LCS-KM under $\ell_{2}$ perturbations with size $\epsilon=128/255$. In this experiment, we augment CIFAR-10 with either external unlabeled data sourced from Tiny-ImageNet or synthetic data generated via DDPM. We consider three configurations of selecting $10\%$, $20\%$, and $100\%$ extra unlabeled data. Our findings are shown in 10a and 10b, revealing that selecting $20\%$ unlabeled data achieves nearly equivalent robust accuracy compared to using the full dataset with $\alpha=100\%$. These results again confirm the efficiency and effectiveness of LCS-KM when adopted in SSAT algorithms.
D-B Low Labeled Data Regime
We further study how the choice of the intermediate model and the amount of labeled data impact the final performance. For all experiments, we use a subset of the CIFAR-10 dataset as labeled data, supplemented with $10\%$ of data selected from a pool of 500500K unlabeled images sourced from Tiny-ImageNet. This additional data is selected using the intermediate model. The amount of unlabeled data remains constant across experiments, while the quantity of labeled data is varied. The intermediate model is trained using self-supervised learning (SSL) with a $5\%$ randomly selected subset of the unlabeled dataset, as long as the labeled data ratio is low. In addition, we include a pre-trained CLIP model [52] as an alternative intermediate model for comparison. The results of these experiments are presented in Table VI. The intermediate model’s quality significantly impacts the final model’s robust accuracy. In low-label regimes (e.g., $1\%$ or $10\%$), models trained using our intermediate model yield comparatively lower robust accuracy than those leveraging the pre-trained CLIP model. However, as the amount of labeled data for the intermediate model increases (e.g., 20% or 100%), the final model results are similar to or better than those of its counterparts based on the CLIP intermediate model. This trend demonstrates the effect of the intermediate model on the final robust accuracy. These observations underscore the importance of having a strong intermediate model for robust performance. Besides, we observe that pre-trained models like CLIP can be a good alternative to training an intermediate model, especially in low-labeled-data regimes. Detailed experiment results and discussions on the pre-trained CLIP intermediate model are provided in Appendix C-B.
D-C Further Discussion
Our work introduces several data reduction strategies that tackle the key challenges in adversarially robust learning, particularly the significant memory and computational overhead of state-of-the-art SSAT methods. By prioritizing boundary-adjacent data, these processes aim to enhance model robustness. However, these methods assume that the model’s latent representations are reliable for clustering and confidence evaluation, an assumption that may falter in the case of undertrained or poorly calibrated models. While these approaches effectively reduce the volume of unlabeled data required for training, their performance can be sensitive to hyperparameter configurations, such as the balance between boundary and non-boundary points and the size of the unlabeled dataset. By focusing on challenging examples near the decision boundary, these selection methods naturally address some aspects of robust overfitting and support robust generalization. Nevertheless, a deeper investigation is required to fully comprehend how these strategies influence decision boundary geometry and the model’s capacity for robust generalization. Combining these selection techniques with fast and efficient adversarial example generation methods, such as fast and free adversarial training, holds promise for further accelerating the training process while preserving robust performance.
Reducing the overall time required for adversarial training is critical for improving the efficiency and scalability of robust learning. Broadly, there are two complementary strategies to achieve this: (i) decreasing the computational cost of each training epoch, and (ii) reducing the total number of epochs needed for convergence. The first strategy aims to make each epoch faster without sacrificing robustness. Several prior works [18, 19, 53, 54] focus on this direction by introducing more efficient attack generation and optimization schemes. These approaches demonstrate that careful algorithmic design can substantially reduce per-epoch runtime, making adversarial training far more practical in large-scale settings. However, as shown in previous studies [15, 13], the total number of training epochs grows sharply with the size of the dataset. Consequently, even if the cost per epoch is reduced, increasing the amount of training data can still lead to a dramatic rise in overall training time. For example, adversarial training with approximately 11M generated samples typically requires around $400$ epochs to converge, whereas training with 100100M samples may demand as many as $2000$ epochs. Developing strategies that mitigate this dependence so that large-scale training does not necessitate a proportional increase in epochs would provide substantial computational savings and further improve the practicality of adversarially robust learning.